OpenStudy (anonymous):

Least To Greatest- 1/2 , 4/3 , 8/8 ,6/5 , 7/4

OpenStudy (anonymous):

let's reduce 8/8 to 1: 1/2, 4/3, 1, 6/5, 7/4 Now what's the lowest common denominator we can find? I think it's 60 (3 * 4 * 5): 1/2 = 30/60 4/3 = 80/60 1 = 60/60 6/5 = 72/60 7/4 = 105/60 Now order from least to greatest.